OpenStudy (anonymous):

Rewrite the rational exponent as a radical. (5^3/4)^2/3

OpenStudy (solomonzelman):

\[(a^b)^c=a^{~b \times c}\]

OpenStudy (solomonzelman):

So you are getting\[5^{~3/4~\times ~2/3}=5^{~1/2}=\sqrt[2]{5^1}=\sqrt{5}\]

OpenStudy (anonymous):

oh okay thank you soo much

OpenStudy (solomonzelman):

` A n y t i m e ! `